OpenStudy (anonymous):

A carton has a length of fraction 2 and 1 over 4 feet, width of fraction 1 and 3 over 5 feet, and height of fraction 2 and 1 over 3 feet. What is the volume of the carton?

OpenStudy (anonymous):

I assume these are mixed fractions? First let's convert these mixed fractions into improper fractions A simple formula to convert a mixed fraction into improper fraction is as follows: \[A \frac{x}{y}=\frac{(y \times A)+x}{y}\] Consider the simple fraction \[2\frac{1}{3}\] This becomes, by our formula \[\frac{2 \times 3+1}{3}=\frac{7}{3}\] If you think about it intuitively Now 2 and a third implies 2.33.... and if you divide 7 by 3 in a calculator you will indeed get 2.33.... After converting into mixed fractions you work is as simple as multiplying the 3 dimensions For a cuboidal object of length l, breadth(or width) b and height h, it's volume V is given by \[V=l \times b \times h=lbh\]

OpenStudy (anonymous):